# Formation and Study of a Spherical Plasma Liner for Plasma-Jet-Driven Magneto-Inertial Fusion

### 20 Jan 2024 v1

@@ IV.1 Imaging
A series of images from the Pixelfly fisheye cameras is shown Fig. 7, and illustrates the general behavior from jet ejection to initial merge and finally to central stagnation. The liner assembly begins between 11 and 17 s. Stagnation and the brightest moment occurs around 46 s. It should be noted that because each image corresponds to a different shot, jet velocities may deviate slightly between each image as a result of shot-to-shot variation (see Fig. 11). In contrast, the set of images shown in Fig. 8 are from a single shot using the multi-frame camera Hadland camera. Both the Pixelfly images and Hadland images indicate a uniform liner is indeed formed by the 36 discrete jets with shockless morphology en route to chamber center, as opposed to earlier observations using a smaller subset of guns [18, 17] operating at different conditions. Hadland multi-frame imaging of the process for a single shot also shows a liner stagnation which maintains good sphericity, deviating slightly in some frames towards a more oblate shape. This may be related to the 9 TL guns on the bottom of the chamber which necessarily fire with slightly different settings than the other guns as discussed in Section II. The dissipation time that the overall emission strength recedes back to background levels is around 40 s after stagnation.
IV.2 Jet Velocities and Liner Merge
@@
Velocities averaged over a set of 30 shots are shown in Fig. 10. Uncertainties are more so a result of shot-to-shot variation rather than in the analysis methodology. This variation is visualized in Fig. 11, in which the three shots shown give rather different velocity values, which can be qualitatively confirmed by noting the separation of the two signals in e.g. subplot a. versus those in b. Although shot-to-shot variation is evident in the magnitude of the error bars, mean values over the shot series are generally fall within 10 km/s of the targeted 50 km/s with few outliers - the tuning of gun 34 for example was ineffective as the reduction in GV voltage necessary to elevate its velocity also prevented its firing at all. The velocity determination method discussed in Section III.2 in about 10% of cases return clearly erroneous velocity values (by-hand calculation based on signal separation in these cases sharply disagrees with those calculated via the cross correlation method) - to discard these as outliers, for a given jet on a given shot, only calculated front velocities between 20 and 85 km/s are accepted. With present gun settings it is unlikely jets are fired at velocities truly outside these bounds.
IV.3 Spatially Resolved Spectroscopy
@@
Performing the comparison to PrismSPECT for all fibers at times throughout liner formation and implosion provides a number of interesting insights, summarized in Fig. 13. There is a center-wards (lower fiber number) decrease in Ar amount observed in subfigure b. is not attributable to changes in electron temperature which remains at about 2.5 0.18 eV over space and time as indicated in subfigure a., and is instead evidence of the residual chamber air being swept up by the liner front. The unchanging nature of was also observed in an earlier PLX study involving merging of only a few guns [18]. Another striking feature of the spectra is evolution of continuum emission intensity from increasing Bremsstrahlung and bound-free emissions illustrated in subfigure c. This provides relative plasma density behavior spatially and temporally, as the continuum intensity scales as , as discussed in Section III.5.
@@ IV.4 High Resolution Spectroscopy
The high resolution spectrometer described in Section III.3 captures a finely resolved profile of the Ar II 480.60 nm emission line. This profile shape originates through a convolution of several mechanisms - the easiest of these to analyze is the instrumental profile whose typically Gaussian shape is the profile constituent originating from optical components, not the plasma itself. Stark broadening manifests as a Lorentzian profile in response to the electric microfield from electrons and ions in the vicinity of emitting ions, and therefore scales linearly with the plasma density. Associated Stark widths for different densities for this line are obtained in Ref. 24. The last profile contribution is a composite effect from each jet conferring a Doppler shifted profile with the previously mentioned broadening freatures due to the projection of the jet’s radial velocity onto the spectral LOS.
To analyze the share of these attributes in the observed emission profile, a technique was developed that models the incoming jets as cylinders of uniform diameter and singular speed rather than a distribution of speeds. These cylindrical jets intersect the instrument’s LOS which itself is approximated as a narrower cylinder at an angle corresponding to the gun port location relative to the collection optics port. As a result, jets originating close to the collection optics port and close to the polar opposite location will contribute a greater proportion of light to the overall profile because they have a greater intersection volume, and will also be Doppler shifted the most due to the shallow angle of jet with respect to optics LOS. Using this framework, resultant composite profiles can be reconstructed for a set of values of plasma bulk velocities and densities, and seeing which best match the observed data, bearing in mind that the simplicity of the analysis model will contribute some uncertainty.
@@ IV.5 Interferometry
Inspection of the interferometry traces on both individual shots and multi-shot averages in Fig. 15 reveals several features. As the leading edge of the jets arrive, a slow increase of density is observed, preceding a sharp increase on the central chord. At later time, similar sharp increases are observed on the off-radial chords. In comparison with the experimental images in Fig. 8, it is clear that the sharp increase is connected to the formation of the expanding luminous region on center, which is the collisional stagnation of the liner and subsequent expansion / rebound of the stagnated liner plasma. The paired chords display decent symmetry within shot-to-shot variation on the early time rise of the density, with variations becoming observable on the rebound in both magnitude and timing of the rebound arrival. This is likely indicative of a small systematic asymmetry in the jet-to-jet delivered mass and pressure, leading to the observed asymmetry in the rebound over many shots. Spatial and temporal behavior of electron density determined from MCS continuum intensity as shown in Fig. 16 are in general agreement with that determined via interferometry, and the time of peak density corresponds well with the bright stagnation time seen in Figs. 7 and 8.
V Discussion:
@@ V.1 Liner Uniformity
A major question in PJMIF is whether a sufficiently smooth and uniform spherical liner can be formed from the merging of discrete hypersonic jets, particularly in the face of density perturbations from shock waves between the merging jets. The inertial fusion field in particular has encountered many manifestations of the Rayleigh-Taylor instability when attempting liner implosions, and found in all cases that smooth liners and shells of the utmost uniformity are essential to avoid enhanced instability growth. In these experiments and has been noted in previous works, the absence of the sharp density gradients caused by shock waves upon initial plasma jet merging can be attributed to significant kinetic interpenetration between the jets [25, 26, 17]. The characteristic ion-ion interpenetration length is given by:
@@
where is the Coulomb logarithm between counterstreaming ions and , is the mean charge state, is the ion/proton mass ratio, is the energy associated with velocity , in units of eV, and is density in cm-3. If ions and are of the same element (argon in this case), then the interpenetration length scales as . As such, the velocity, merging angle, and charge state of the jets exert a large control over the degree of interpenetration. If plasma-jet-driven liner implosions are scaled up to greatly increased plasma density, one can expect decreased jet interpenetration and re-emergence of collisional shock waves between the primary plasma jets. Due to the strong dependence of interpenetration length on velocity, increases in jet velocity concomitant with increases in liner density would have offsetting effects, resulting in the persistence of significant interpenetration even at many orders of magnitude greater energy densities. The exact termination point of this scaling is not empirically known, but we conjecture that it will at maximum end with densities corresponding to the onset of optical thickness of the liner plasma, inhibiting effective radiative cooling of the liner and lead to an increase in liner average charge state , leading to the rapid onset of liner collisionality. Future work may shed light on these and other detailed aspects of the jet merging and convergence phase and how they may be affected when scaling towards fusion-relevant conditions. The current experiment provides effectively, a single encouraging data point, showing that a liner free of primary shock waves can be assembled at these conditions.
V.2 Liner Ram Pressure
@@
We can determine bounds of peak ram pressure () by combining the initial jet velocities measured from photodiodes () for the upper bound velocity, and high-resolution () spectroscopy for lower bound velocity, plasma density from interferometry, and from MCS results to obtain . To estimate from interferometry we take the peak line-integrated value of about cm-2, over the apparent domain that imaging indicates to be brightest - around 15 cm. Thus the upper bound of ram pressure is about 123 bar, and the lower bound is about 20 bar. This level is consistent with modeling predictions that have incorporated the present jet lengths of cm and moderate Mach numbers , revealing that there is a significant loss in liner density and ram pressure incurred due to the volumetric expansion of the individual jets en route to the liner merging radius, as well the extended length of the individual jets. If a thinner liner with sharper radial density profile can be formed, possibly by use of jets with higher Mach number and optimized launch velocity profile, the ram pressure has potential to be increased by many orders of magnitude at a given energy scale. Another path to increasing liner ram pressure and energy density is to simply decrease the liner standoff distance, effectively making the chamber smaller and starting the jets at a lesser radius. While this is attractive for the liner performance, it puts increased demands on the eventual system components to survive in a harsher fusion heat and radiation environment, and may complicate reactor engineering [27].
VI Conclusions:
@@
The present set of experiments at PLX delivers the most substantial experimental measurements to date to investigate the feasibility of the PJMIF liner formation approach. Spatial and temporal measurements are obtained of an argon liner formed via the merger of 36 discrete plasma jets, following the liner evolution through jet launch, merging, spherical convergence, stagnation, and rebound. Notably, conditions of moderate interpenetration between primary jets are observed to lead to the absence of dense primary shock structures in the early liner, allowing increased density uniformity to be achieved. If primary shock waves are mitigated in this way, it seems that arbitrarily smooth liners may be able to be formed with corresponding improvements in primary jet parameters and control. A clear transition from initial kinetic interpenetration to collisional interaction is observed, in accordance with theoretical expectation. The obtained diagnostic measurements also provide achieved absolute parameter values of ram pressure, stagnation density and temperature, useful for benchmarking computational models and understanding the scaling of plasma liners to increased energy density. The results provide a first singular data point for these benchmarking purposes – future experiments could seek to verify theoretical liner scalings by exploring the effect of perturbations to liner parameters.
@@ Acknowledgements.
The authors would like to thank Tom Byvank, Edward Cruz, Levi Grantz, Ricardo Martinez, and many others in the community who have contributed greatly to the PLX project and the current experiments, as well as the pioneers of the PJMIF concept and PLX experiment, Francis Thio and Scott Hsu, respectively. The information, data or work presented herein was funded in part by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy under Award Numbers 20/CJ000/03/02 and DE-AR0001268, BETHE program. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof. This work was supported by the U.S. Department of Energy in part through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).Data Availability

### 20 Feb 2024 v2

@@ IV.1 Imaging
A series of images from the Pixelfly fisheye cameras is shown Fig. 7, and illustrates the general behavior from jet ejection to initial merge and finally to central stagnation. The liner assembly begins between 11 and 17 s. Stagnation and the brightest moment occurs around 46 s. It should be noted that because each image corresponds to a different shot, jet velocities may deviate slightly between each image as a result of shot-to-shot variation (see Fig. 11). In contrast, the set of images shown in Fig. 8 are from a single shot using the multi-frame Hadland camera. Both the Pixelfly images and Hadland images indicate a uniform liner is indeed formed by the 36 discrete jets with shockless morphology en route to chamber center, as opposed to earlier observations using a smaller subset of guns [21, 20] operating at different conditions. Hadland multi-frame imaging of the process for a single shot also shows a liner stagnation which maintains good sphericity, deviating slightly in some frames towards a more oblate shape. This may be related to the 9 TL guns on the bottom of the chamber which necessarily fire with slightly different settings than the other guns as discussed in Section II. The dissipation time that the overall emission strength recedes back to background levels is around 40 s after stagnation.
IV.2 Jet Velocities and Liner Merging
@@
Velocities averaged over a set of 30 shots are shown in Fig. 10. Uncertainties are more so a result of shot-to-shot variation rather than in the analysis methodology. This variation is visualized in Fig. 11, in which the three shots shown give rather different velocity values, which can be qualitatively confirmed by noting the separation of the two signals in e.g. subplot a. versus those in b. Although shot-to-shot variation is evident in the magnitude of the error bars, mean values over the shot series generally fall within 10 km/s of the targeted 50 km/s with few outliers - the tuning of gun 34 for example was ineffective as the reduction in GV voltage necessary to elevate its velocity also prevented its firing at all. The velocity determination method discussed in Section III.2 returns clearly erroneous velocity values in about 10% of cases (by-hand calculation based on signal separation in these cases sharply disagrees with those calculated via the cross correlation method). To discard these as outliers, for a given jet on a given shot only calculated front velocities between 20 and 85 km/s are accepted. With present gun settings it is unlikely jets are fired at velocities truly outside these bounds.
IV.3 Spatially-Resolved Emission Spectroscopy
@@
Performing the comparison to PrismSPECT for all fibers at times throughout liner formation and implosion provides a number of interesting insights, summarized in Fig. 13. There is a center-wards (lower fiber number) decrease in Ar amount, observed in subfigure b. This decrease is not attributable to changes in electron temperature which remains at about 2.5 0.18 eV over space and time as indicated in subfigure a., and is instead evidence of the residual chamber air being swept up by the liner front. The unchanging nature of was also observed in an earlier PLX study involving merging of only a few guns [21]. Another striking feature of the spectra is evolution of continuum emission intensity from increasing Bremsstrahlung and bound-free emissions illustrated in subfigure c. This provides relative plasma density behavior spatially and temporally, as the continuum intensity scales as , as discussed in Section III.5.
@@ IV.4 High Resolution Spectroscopy
The high resolution spectrometer described in Section III.3 captures a finely resolved profile of the Ar II 480.60 nm emission line. This profile shape originates through a convolution of several mechanisms - the easiest of these to analyze is the instrumental profile whose typically Gaussian shape is the profile constituent originating from optical components, not the plasma itself. Stark broadening manifests as a Lorentzian profile in response to the electric microfield from electrons and ions in the vicinity of emitting ions, and therefore scales linearly with the plasma density. Associated Stark widths for different densities for this line are obtained in Ref. 30. The last profile contribution is a composite effect from each jet conferring a Doppler shifted profile with the previously mentioned broadening freatures due to the projection of the jet’s radial velocity onto the spectral LOS.
To analyze the share of these attributes in the observed emission profile, a technique was developed that models the incoming jets as cylinders of uniform diameter and singular speed rather than a distribution of speeds. These cylindrical jets intersect the instrument’s LOS which itself is approximated as a narrower cylinder at an angle corresponding to the gun port location relative to the collection optics port. As a result, jets originating close to the collection optics port and close to the polar opposite location will contribute a greater proportion of light to the overall profile because they have a greater intersection volume, and will also be Doppler shifted the most due to the shallow angle of jet with respect to optics LOS. Using this framework, resultant composite profiles can be reconstructed for a set of values of plasma bulk velocities and densities and seeing which best match the observed data (bearing in mind that the simplicity of the analysis model will contribute some uncertainty).
@@ IV.5 Interferometry
Inspection of the interferometry traces on both individual shots and multi-shot averages in Fig. 15 reveals several features. As the leading edge of the jets arrive, a slow increase of density is observed, preceding a sharp increase on the central chord. At later time, similar sharp increases are observed on the off-radial chords. In comparison with the experimental images in Fig. 8, it is clear that the sharp increase is connected to the formation of the central expanding luminous region, which is the collisional stagnation of the liner and subsequent expansion / rebound of the stagnated liner plasma. The paired chords display decent symmetry within shot-to-shot variation on the early time rise of the density, with variations becoming observable on the rebound in both magnitude and timing of the rebound arrival. This is likely indicative of a small systematic asymmetry in the jet-to-jet delivered mass and pressure, leading to the observed asymmetry in the rebound over many shots. Spatial and temporal behavior of electron density determined from MCS continuum intensity, as shown in Fig. 16, are in general agreement with that determined via interferometry, and the time of peak density corresponds well with the bright stagnation time seen in Figs. 7 and 8.
V Discussion
@@ V.1 Liner Uniformity
A major question in PJMIF is whether a sufficiently smooth and uniform spherical liner can be formed from the merging of discrete hypersonic jets, without significant density perturbations from shock waves between the merging jets. The inertial fusion field in particular has encountered many manifestations of the Rayleigh-Taylor instability when attempting liner implosions, and found in all cases that smooth liners and shells of the utmost uniformity are essential to avoid enhanced instability growth.
In previous PLX studies of obliquely merging jets in pairs, trios, and clusters of 6 and 7 PLX plasma guns, inter-jet shocks along the merge plane of different jets were frequently observed, whereas it appears largely absent in the present experiment. This key difference is consistent with the former experiments typically exhibiting slower velocities and shallower mean merge angles between adjacent jets. In Ref[20] it was found that when gun angular separations were increased from 11.6 to 20.5 degrees half-angle, interpentration length exceeded ion mean free path length and thus shock formation was substantially mitigated. In Ref[21] significant improvements to jet-jet mass uniformity and modest improvements to velocity uniformity were implemented. Measurements of the merger of clusters of 6 or 7 jets indicated increased temperature and ionization, but could potentially still remain in the desired interpenetrating regime without shock formation.
In the present experiment, the mean separation angle between adjacent jets of 36 degrees constitutes the primary reason for not observing the inter-jet shocks observed in prior experiments, since their merge velocity is greater for larger angles. The characteristic ion-ion interpenetration length is given by:
@@
where is the Coulomb logarithm between counterstreaming ions and , is the mean charge state, is the ion/proton mass ratio, is the energy associated with velocity , in units of eV, and is density in cm-3. If ions and are of the same element (argon in this case), then the interpenetration length scales as . As such, the velocity, merging angle, and charge state of the jets exert a large control over the degree of interpenetration.
@@
If plasma-jet-driven liner implosions are scaled up to greatly increased plasma density, one can expect decreased jet interpenetration and re-emergence of collisional shock waves between the primary plasma jets. Due to the strong dependence of interpenetration length on velocity, increases in jet velocity concomitant with increases in liner density would have offsetting effects, resulting in the persistence of significant interpenetration even at many orders of magnitude greater energy densities. The exact termination point of this scaling is not empirically known, but we conjecture that it will at maximum end with densities corresponding to the onset of optical thickness of the liner plasma, inhibiting effective radiative cooling of the liner and leading to an increase in liner average charge state . Ultimately, this chain of events may lead to a relatively rapid onset of the liner collisionality occurring at a threshold in parameter space. Future work may shed light on these and other detailed aspects of the jet merging and convergence phase and how they may be affected when scaling towards fusion-relevant conditions. The current experiment effectively provides a single encouraging data point showing that a liner free of primary shock waves can be assembled at these conditions.
V.2 Comparison to Hydrodynamic Simulation
The experimental campaign is compared to smooth-particle hydrodynamics (SPH) simulations performed with the SPFMax code at UAH [31]. Simulations to date have been performed with uniform nominal jet initial conditions of density and velocity profile, not accounting for individual or shot-to-shot variations between jets. No effects of any residual magnetic field are included in the simulations (consistent with the cold dense jets being largely unmagnetized post-flight). An ion-slip model is used to account for the effects of kinetic interpenetration, leveraging the particle-based mechanics of the SPH code – collisional interaction between fluid particles is adjusted by a multiplying factor based on equation 6, with the effects that particles of high relative velocity experience little collisional interaction, and vice versa. Results of stagnation density profile with the ion-slip model turned on are shown in Figure 17, showing similar morphology to the experimental imaging and inferred density profiles. It is observed that simulation results with ion slip model turned on agree much more closely with the experimental results, re-enforcing the view that interpenetration is playing a strong role in high velocity liner formation.
V.3 Liner Ram Pressure
@@
We can infer bounds of liner peak ram pressure () by combining the initial jet velocities measured from photodiodes () for the upper bound velocity, and high-resolution () spectroscopy for lower bound velocity, plasma density from interferometry, and from MCS results to obtain . To estimate from interferometry we take the peak line-integrated value of about cm-2, over the apparent domain that imaging indicates to be brightest - around 15 cm. Thus the upper bound of ram pressure is about 123 bar, and the lower bound is about 20 bar. This level is consistent with modeling predictions that have incorporated the present jet lengths of cm and moderate Mach numbers , revealing that there is a significant loss in liner density and ram pressure incurred due to the volumetric expansion of the individual jets en route to the liner merging radius, as well the extended length of the individual jets. If a thinner liner with sharper radial density profile can be formed, possibly by use of jets with higher Mach number and optimized launch velocity profile, the ram pressure has potential to be increased by many orders of magnitude at a given energy scale. Another path to increasing liner ram pressure and energy density is to simply decrease the liner standoff distance, effectively making the chamber smaller and starting the jets at a lesser radius. While this is attractive for the liner performance, it puts increased demands on the eventual system components to survive in a harsher fusion heat and radiation environment, and may complicate reactor engineering [32].
VI Conclusions
@@
The present set of experiments at PLX delivers the most substantial experimental measurements to date to investigate the feasibility of the PJMIF liner formation approach. Spatial and temporal measurements are obtained of an argon liner formed via the merger of 36 discrete plasma jets, following the liner evolution through jet launch, merging, spherical convergence, stagnation, and rebound. Notably, conditions of moderate interpenetration between primary jets are observed to lead to the absence of dense primary shock structures in the early liner, allowing increased density uniformity to be achieved. If primary shock waves are mitigated in this way, it seems that arbitrarily smooth liners may be able to be formed with corresponding improvements in primary jet parameters and control. A clear transition from initial kinetic interpenetration to collisional interaction is observed, in accordance with theoretical expectation. The obtained diagnostic measurements also provide achieved absolute parameter values of ram pressure, stagnation density, and temperature, useful for benchmarking computational models and understanding the scaling of plasma liners to increased energy density. The results provide a first singular data point for these benchmarking purposes – future experiments could seek to verify theoretical liner scalings by exploring the effect of perturbations to liner parameters.
@@ Acknowledgements.
The authors would particularly like to thank Tom Byvank, Edward Cruz, Levi Grantz, Ricardo Martinez, and many others who have contributed greatly to the current experiments and overall PLX project, as well as the pioneers of the PJMIF concept and PLX experiment, Francis Thio and Scott Hsu, respectively. The information, data or work presented herein was funded in part by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy under Award Numbers 20/CJ000/03/02 and DE-AR0001268, BETHE program. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof. This work was supported by the U.S. Department of Energy in part through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).Data Availability